OpenStudy (redohawk):

given tan0=4/3 and sin(-4/5) find cos 0

OpenStudy (zarkon):

I assume you mean given \(\tan(\theta)=\frac{4}{3}\) and \(\sin(\theta)=\frac{-4}{5}\) find \(cos(\theta)\)

OpenStudy (redohawk):

yes

OpenStudy (campbell_st):

so you need a quadrant where tan is positive and sin is negative you have also been given the 3 sides of a right triangle

OpenStudy (redohawk):

A. 1/3 B. -1/3 C. 3/5 D. -3/5

OpenStudy (redohawk):

@.Sam.

OpenStudy (redohawk):

@SnuggieLad

OpenStudy (zarkon):

\[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\] \[\cos(\theta)=\frac{\sin(\theta)}{\tan(\theta)}\]

OpenStudy (dumbcow):

Set up a triangle using the given ratios sin = opp/hyp tan = opp/adj  cos = adj/hyp = -3/5